A sample of neon gas diffuses through a porous barrier at a rate of 48.7 grams per hour (g/hr). At what rate (in the same units) does argon gas diffuse through the same barrier if the temperature is the same?

Enter your answer in decimal form with the correct number of sig figs.
Answer #1

By Graham's law of Diffusion, the rate of Effusion is inversely proportional to square if it's mass

R1/R2= sq.root (M2/M1)

48.7/R2= sq.root(40/20)

48.7/R2= 1.414

R2= 48.7/1.414= 34.4 g/hr